OpenStudy (anonymous):

Direct Comparison Test/ Limit Comparison Tests Use one of these tests to find if they are divergent or convergent I can actually find by integrating these improper integrals but I'm supposed to use one of these tests. It would be great if the limit comparison test was used because I don't know how to choose the function for the comparison. Explanations please! Questions:

OpenStudy (amistre64):

the choice of function is more intuitive than anything

OpenStudy (anonymous):

\[\int\limits\limits_{\pi}^{\infty} (1-\sin(x))/x^2\] \[\int\limits_{2}^{\infty} xdx/\sqrt{x^4-1}\]

OpenStudy (anonymous):

I chose \[x/\sqrt{x^4} \] and \[1/\sqrt{x^4}\] but it ended up getting simplfied and ended up a hard to solve integral or ended up with divergent which is not true. The first one, I have zero idea on what to choose. There are other problems but I hope if I understand this, I can solve them by myself.

OpenStudy (amistre64):

\[\frac{1-sin(n)}{n^2}\] for large values of n that might resemble \(\frac{-sin(n)}{n^2}\) which really looks to mimic \(-\frac{1}{n^2}\) but it might be good to get a view of it on the wolf first

OpenStudy (anonymous):

It might sound rude but I really don't understand the point of these tests when it might be simpler to just integrate them. Arrrgh! >:(

OpenStudy (amistre64):

the point is just for gaining a bit of experience .... some hands on work if you will :)

OpenStudy (amistre64):

im not to adept at these comparisons either ...

OpenStudy (anonymous):

:/